OpenStudy (anonymous):

find the equation of a sphere that has endpoints of a diameter at A( 1,2,3) and B(-1,0,6) midpoint of line AB = (0,1,9/2) length = root(17)

OpenStudy (anonymous):

i THINK its \[x^2 + (y-1)^2+ (z-9/2)^2 = \sqrt(17)/2\] amirite?

Directrix (directrix):

(x,y,z) is on the sphere with center (h,j,k) and radius r if and only if (x - h)^2 + (y - j)^2 + (z - k)^2 = r^2 That's what I got. Same center and same radius.

OpenStudy (anonymous):

K ^.^ thank you :D